# Functional Near-Infrared Spectroscopy (fNIRS) Analysis of Interaction Techniques in Touchscreen-Based Educational Gaming

### v5

@@ 7. Conclusion
In conclusion, this study evaluated hand and stylus usage in an in-house educational game, measuring oxygenated hemoglobin concentration ($\Delta HbO$), relative neural efficiency, and involvement alongside feedback from user experience measures (SUS and NASA-TLX). The findings indicate a statistically significant difference in $\Delta HbO$, relative neural efficiency and involvement between the two input methods. A significant difference was shown in user experience metrics, as evidenced by SUS and NASA-TLX scores. In particular, relative neural efficiency suggests that the stylus requires more motor skill and coordination, potentially increasing neural involvement through higher oxygen flow compared to using the hand, which is more natural and intuitive. These findings highlight the importance of accounting for cognitive and physiological factors in educational games and natural user interfaces. Understanding how input modalities influence user performance and experience can help developers optimize learning outcomes. This study provides valuable insights for designers seeking to choose input modalities that enhance user performance and learning in educational games.
Safe and Responsible Innovation Statement This study was approved by the Institutional Review Board and conducted with strict safeguards to ensure participant anonymity and privacy. We examined how different input modalities affect cognitive effort in educational gameplay using non-invasive fNIRS signals. Data collection and analysis followed ethical protocols, and participant diversity was considered to minimize bias. We advocate for transparent reporting, ethical handling of neural data, and responsible application of brain-sensing insights in educational contexts.
Acknowledgement
We express our gratitude to the study participants and lab members. We also thank the National Science Foundation for its support.Any opinions, findings, and conclusions expressed in this material are those of the authors and do not reflect the views of sponsors.

### v6

@@ 7. Conclusion
In conclusion, this study evaluated hand and stylus usage in an in-house educational game, measuring oxygenated hemoglobin concentration ($\Delta HbO$), relative neural efficiency, and involvement alongside feedback from user experience measures. The findings indicate a statistically significant difference in $\Delta HbO$, relative neural efficiency and involvement between the two input methods. A significant difference was shown in user experience metrics, as evidenced by SUS and NASA-TLX scores. In particular, relative neural efficiency suggests that the stylus requires more motor skill and coordination, potentially increasing neural involvement through higher oxygen flow compared to using the hand, which is more natural and intuitive. These findings highlight the importance of accounting for cognitive and physiological factors in educational games and natural user interfaces. Understanding how input modalities influence user performance and experience can help developers optimize learning outcomes. This study provides valuable insights for designers seeking to choose input modalities that enhance user performance and learning in educational games.
Safe and Responsible Innovation Statement The study was approved by the IRB and conducted with strict privacy safeguards. We used non-invasive fNIRS to examine how input modalities affect cognitive effort in educational gameplay. Data collection followed ethical protocols, with attention to participant diversity. We promote transparency and responsible use of neural data in educational context.
Acknowledgment
We express our gratitude to the study participants and lab members. We also thank the National Science Foundation for its support (#$2222661-2222663$, #$2321274$, and #$2426003$). Any opinions, findings, and conclusions expressed in this material are those of the authors and do not reflect the views of sponsors.